OpenStudy (anonymous):

How do I solve: (4x-1)(3x+7) = (5x-1)(2x+3)-6? The answer is supposed to be: -3+/-2√2.. how?

OpenStudy (anonymous):

I got \[72x ^{2} + 103x -25 =0\] and do I use quadratic formula?

OpenStudy (dumbcow):

i get: 2x^2 +12x+2 = 0

OpenStudy (nenadmatematika):

something's wrong with your calculation...you can't get 72x^2...:D

OpenStudy (anonymous):

really?

OpenStudy (dumbcow):

but yes you would use quadratic formula

OpenStudy (anonymous):

hooow>

OpenStudy (dumbcow):

(4x-1)(3x+7) = 12x^2 -3x +28x -7 (5x-1)(2x+3)-6 = 10x^2 -2x +15x -3 -6

OpenStudy (anonymous):

for the (5x-1)(2x+3)-6, i don't multiply -6 to the rest of the terms?

OpenStudy (dumbcow):

only if its (-6)*(5x-1)(2x+3) the way its written it implies subtracting 6

OpenStudy (anonymous):

really? ohh.. o.o thank you!!

OpenStudy (dumbcow):

yw